Hello, I need help with this Statement/Reasons type question.
Complete the following solution by filling in the missing statements and reasons.
•Statement: x-2(3x-4)=8 Reason:Given
•Statement:_________ Reason:Distributive Property
•Statement:-5x+8=18 Reason:Simplify
•Statement:_________ Reason:____________
•Statement: x=-2 Reason:____________
Thank you so much (:

Answer by jim_thompson5910(35256)


The filled in answers are given in red
NumberStatementReason
1x-2(3x-4)=18Given
2x-6x+8=18Distributive Property
3-5x+8=18Simplify
4-5x=10Subtraction Property of Equality
5x = -2Division Property of Equality


Notes:
  • In step 2, we multiply the -2 by each term inside the parentheis.
  • In step 4, we subtract 8 from both sides.
  • In step 5, we divide both sides by -5 to isolate x